Find the marginal cost functions for each of the following average cost functions.




a) AC = 1.5Q + 4 + 46




𝑄




b) AC =




160




𝑄




+ 5 - 3Q + 2Q2

1
Expert's answer
2022-03-14T13:16:59-0400

"MC=\frac{dTC}{dQ}"

Where "TC=AC\u00d7Q"

a) "AC=1.5Q+\frac{4}{Q}+46"

"TC=(1.5Q+\frac{4}{Q}+46)Q"

"TC=1.5Q^2+4+46Q"

"MC=3Q+46"


b) "AC=\frac{160}{Q}+5-3Q+2Q^2"

"TC=160+5Q-3Q^2+2Q^3"

"MC=5-6Q+6Q^2"